Bramble Company purchased a new plant asset on April 1, 2020, at a cost of $753,000. It was estimated to have a service lifePlease also do for double declining balance! Thanks!
Answer #1

Sum of Year's Digits Method

>> Sum of years = 20 ( 20 + 1 ) / 2 = 210

>> Depreciation for 2020 = ( $ 753,000 - $ 64,200 ) * ( 20 / 210 ) * ( 9 / 12 )

>> Depreciation for 2020 = $ 49,200.

>> Remaining period = 20 - ( 9 /12 ) = 19.25.

>> Depreciation of 2021 = ( $ 753,000 - $ 64,200 ) * ( 19.25 / 210 )

>> Depreciation of 2021 = $ 63,140.

___________________________________________________________________

Double declining balance method.

>> Depreciation rate = 200 / 20 = 10 %.

>> Depreciation for 2020 = $ 753,000 * 10 % * ( 9 / 12 ) = $ 56,475.

>> Depreciation for 2021 = ( $ 753,000 - $ 56,475 ) * 10 % = 69,653

____________________________________________________________________